# A MIND for Reasoning: Meta-learning for In-context Deduction

### v1

@@ 4 Experimental Setup
4.1 Models
We run experiments using the Qwen 2.5 family of decoder-only LMs (Qwen Team, 2025). More specifically, we test three sizes: 1.5B, 3B, and 7B parameters. This family of models is selected because it allows us to experiment with varying small sizes (ranging from 1.5 to 7 billion parameters) and achieves a better size vs. performance trade-off compared to other open weights model families.
In addition to the Qwen 2.5 family, we also evaluate the closed-source LLM GPT-4o (OpenAI, 2024) and the Large Reasoning Model (LRM) o3-mini (OpenAI, 2025) on the logical premise selection task. [^8] Note that LRMs are also LLMs, but post-trained to generate longer intermediate chains of thought, improving performance on complex reasoning tasks (Xu et al., 2025). We conduct the evaluation both in a zero-shot setting and in a few-shot setting, using the $S^{\mathrm{supp}}$ study pairs as examples. [^9] See the API details and the exact prompts used to evaluate closed models in Appendix C.3.
@@ 4.1 Models
Fine-tuning Qwen-2.5 1.5B MIND 93.11 ± 0.61 94.28 ± 0.61 91.76 ± 0.27
Baseline 85.56 ± 1.24 91.42 ± 0.82 80.56 ± 1.78
Qwen-2.5 3B MIND 96.16 ± 0.44 96.24 ± 0.56 95.55 ± 0.43
Baseline 93.03 ± 1.15 95.34 ± 1.18 90.92 ± 1.27
Qwen-2.5 7B MIND 98.13 ± 0.98 98.26 ± 0.82 97.69 ± 1.40
Baseline 95.76 ± 1.10 97.27 ± 1.22 94.13 ± 0.90
Prompting GPT-4o Few-shot 39.76 52.91 33.51
@@
Table 1: Core generalization. Accuracy (mean ± std) on test inferences across all type-length combinations (All), plus breakdown into the five shortest (Short) and longest (Long) inferences for each of the seven types of inference. Fine-tuned Qwen models use MIND vs. Baseline; GPT-4o and o3-mini use few-shot vs. zero-shot prompting.
4.2 Experiments
We design experiments to evaluate the ability of MIND to teach pretrained small LMs to systematically apply inferences to new, unseen sets of premises —that is, to reason in a formal way by recognizing and instantiating the same underlying structure independently of the $\mathcal{KB}$s’ content.
To ensure consistency, both MIND and the baseline receive inputs at test time in the same format as during training. MIND models are provided as context $(\mathcal{KB},S^{\mathrm{supp}},x^{\mathrm{query}})$, and are tasked to generate $y^{\mathrm{query}}$, while the baseline receives $(\mathcal{KB},x^{\mathrm{query}})$.
@@ 4.2 Experiments
We then consider two more challenging generalizations involving inferences of unseen length. As illustrated in Figure 3, we examine the case of generalizing to longer inferences when the model has only learned from shorter ones (as studied in Saparov et al. 2023), and vice versa —generalizing to shorter inferences after seeing only longer ones. In the logic literature, they are respectively known as recursiveness and compositionality (Vargas Guzmán et al., 2024). To test this, we train exclusively on inferences whose lengths $x$ are $\sigma(t)\leq x\leq\mu(t)-5$, and test on the five longest inferences for each type, i.e., those whose length is $\mu(t)-5<x\leq\mu(t)$. In the second case, we train on inferences with length $\sigma(t)+5\leq x\leq\mu(t)$, and test only on the five shortest inference lengths for each type, i.e., those with length $\sigma(t)\leq x<\sigma(t)+5$. An intuitive representation of these generalizations is provided in Figure 3. Notably, within the MIND approach, we consider two varying types of study examples $S^{\mathrm{supp}}$: the aligned and disaligned sets of study examples, in which each $(x^{\mathrm{supp}},y^{\mathrm{supp}})$ either falls within or outside the range of inference lengths used for testing, respectively. [^10] More precisely, the meanings of aligned and disaligned depend on whether we are evaluating models on unseen shorter or longer inferences. For longer inferences, disaligned includes inferences with lengths $\sigma(t)\leq x\leq\mu(t)-5$, and aligned includes those with lengths $\mu(t)-5<x\leq\mu(t)$. For shorter ones, instead, aligned includes inferences with lengths $\sigma(t)\leq x<\sigma(t)+5$, and disaligned includes those with lengths $\sigma(t)+5\leq x\leq\mu(t)$.
Figure 6, in the Appendix, shows all inference type-length combinations within training and test split in the core and in the length generalization settings. These datasets contain 1,000 and 100 datapoints for each training and testing type–length combination, respectively. To further investigate the performance of MIND in a limited data regime, we also consider the case where only 100 datapoints are available for each training type–length combination.
@@ 4.3 Prediction Accuracy
We consider a model prediction to be correct if the set of premises extracted from the generated text matches the ground truth set of minimal premises. Using this criterion, we measure accuracy both in aggregate, i.e., across an entire test set, and decomposed by each test type-length combination. All models (1.5B, 3B, and 7B) are fine-tuned three times and with different random seeds, thus we report mean and standard deviation of each accuracy.
@@ 5 Results
Qwen-2.5 1.5B MIND 76.42 ± 2.95 91.75 ± 1.10 70.94 ± 2.27 71.13 ± 1.83
Baseline 63.53 ± 1.16 63.53 ± 1.16 56.67 ± 1.22 56.67 ± 1.22
Qwen-2.5 3B MIND 87.61 ± 1.97 95.86 ± 0.70 77.19 ± 3.53 78.53 ± 1.71
Baseline 76.78 ± 1.63 76.78 ± 1.63 71.88 ± 1.49 71.88 ± 1.49
Qwen-2.5 7B MIND 90.03 ± 1.09 96.84 ± 0.15 76.23 ± 2.91 83.41 ± 1.63
Baseline 80.76 ± 2.65 80.76 ± 2.65 71.08 ± 1.55 71.08 ± 1.55
Table 2: Generalization to unseen lengths. Accuracy (mean ± std) of meta-learning and baseline models when trained on short inferences and tested on longer ones or vice versa. In both cases, we compare the settings in which the inferences in the study examples either falls within (Aligned) or outside (Disaligned) the range of inference lengths used for testing. Baseline models have no study examples, hence such difference does not hold for them.
@@ 5.1 Core Generalization
Across all Qwen-2.5 model sizes (1.5B, 3B, 7B), the meta-learning approach consistently yields higher accuracy than the baseline. Performance improves with model scale in both approaches. For example, MIND accuracy increases from 93.11% (1.5B) to 98.13% (7B) on all type-length combinations, with accuracy on shortest inferences rising from 94.28% to 98.26%, and on the longest ones increasing from 91.76% to 97.69%. In contrast, baseline performance rises more slowly —from 85.56% (1.5B) to 95.76% (7B) —and shows a wider drop on the longest inferences, falling as low as 80.56% for the smallest model. Notably, the performance gap between MIND and the baseline narrows as model size increases, suggesting that larger models achieve better core generalization even without meta-learning. It is worth noting that with limited data, MIND’s advantage over the baseline becomes much wider at all sizes, as shown in Appendix D.3.
The closed-source models GPT-4o and o3-mini still underperform compared to Qwen-2.5 models fine-tuned with MIND. Both models perform poorly in the zero-shot setting but improve with few-shot prompting: GPT-4o reaches 39.76% on all type-length combinations (with 52.91% on shortest and 33.51% on longest inferences), while o3-mini performs substantially better (88.45% all combination, 87.91% on shorters, and 88.51% on longest). As expected, performance on the longest inferences is worse than that on the shortest ones for GPT-4o, while o3-mini maintains a more robust performance across inference lengths.
@@ 5.2 Length Generalization
Table 2 shows that MIND models consistently outperform baseline models in generalizing to both longer and shorter inferences than those seen during training. In core generalization, we observed that longer inferences are more challenging than shorter ones. Instead, in the case of unseen lengths, an interesting and somewhat counterintuitive pattern emerges: it is generally easier for models to generalize to longer inferences than to shorter ones. This is true across all model sizes and in both approaches; For instance, the largest model, Qwen-2.5 7B, achieved 90.03% accuracy on longer inferences (disaligned) compared to 76.23% on shorter ones (disaligned).
Aligning study example lengths with the test condition (aligned) proves moderately to highly effective for unseen short inferences across all MIND model sizes. For example, Qwen-2.5 1.5B improved from 76.42% to 91.75%, and Qwen-2.5 3B improved from 87.61% to 95.86%. For unseen long inferences, this alignment is moderately effective in larger models: Qwen-2.5 7B improved from 76.23% to 83.41%, while the 1.5B and 3B models showed smaller gains (70.94% to 71.13% and 77.19% to 78.53%, respectively). These results indicate that MIND enables models in the aligned condition to exploit abstract patterns in the study examples (unseen inference lengths), allowing them to more effectively answer query hypotheses requiring length generalization.
Again, MIND’s better performance in length generalization is especially noticeable with limited training data, where the difference between MIND and baseline models grows significantly (see Appendix D.3 for more details).
@@
L $\rightarrow$ S MIND (aligned) 42.94 4.9 36.68 2.1 57.5
MIND (disaligned) 28.31 3.72 52.81 1.76 66.06
Baseline 28.21 6.19 23.38 2.1 72.78
S $\rightarrow$ L MIND (aligned) 9.76 1.66 87.54 5.08 60.94
MIND (disaligned) 14.14 6.14 81.82 3.65 35.35
Baseline 3.87 2.36 89.79 6.66 66.9
Table 3: Error analysis. Error analysis comparing MIND and baseline on long to short (L $\rightarrow$ S) and short to long (S $\rightarrow$ L) generalization. The table shows percentages and averages for non-minimal valid sets of premises (NVM) and missing necessary $A$ premises (MAP), and the percentage of hallucinated premises (HP).
@@ 6 Error Analysis
Beyond simply measuring the accuracy of MIND and the baseline, we additionally focus on two main types of errors models make when evaluated on unseen lengths. First, among all errors, we consider the proportion of non-minimal valid set of premises (NVM). This means that the correct minimal set was generated by the model, but together with unnecessary premises; for this case, we also measure how many unnecessary premises, on average, the models generate. Alternatively, models may fail to provide the complete $A$-chain within the correct minimal set of premises, meaning that at least one necessary $A$ premise is missing (MAP); here, we also track the average number of missing necessary $A$-formulas in erroneous answers. NVM and MAP are mutually exclusive. Furthermore, we consider an additional type of error that can occur simultaneously with either NVM or MAP: models may hallucinate premises —referred to as hallucinated premises (HP) —and output a formula that is not contained in the $\mathcal{KB}$.
Table 3 presents the error analysis for Qwen-2.5 7B [^11] Each model was fine-tuned three times with different random seeds, we selected the best model for each approach for this analysis. on the challenging length generalization settings. [^12] See Appendix D.4 for further error analysis results. HP is a common error type across both settings (often $>$50%). The baseline model has the highest HP rate in long to short (72.78%), while MIND models are generally better.
When generalizing to shorter inferences, a substantial portion of errors (28-43%) are NVM, indicating models indeed find logical solutions but include unnecessary premises. In this context, a lower number of unnecessary premises is better, as it is closer to the minimal set. The baseline model adds the most unnecessary premises (6.19 average), compared to MIND (disaligned) (4.9) and MIND (aligned) (3.72).
For generalizations to longer inferences, errors show different patterns, with few NVM errors (4-14%) and predominantly MAP errors (81-90%). The average number of missing premises is higher in short to long (3.65-6.66) than in long to short (1.76-2.1), suggesting models struggle to provide the complete set of premises when evaluated on longer inferences than seen during training. The baseline model struggles most with longer inferences, with a high MAP error rate (89.79%) and a large number of missing premises (6.66) contributing to its lower accuracy compared to MIND.
@@ 7.1 LLMs’ Logical Capabilities
Recent work has highlighted weaknesses in LLMs’ logical reasoning. LLMs often struggle with OOD generalization (Clark et al., 2021; Saparov et al., 2023; Vargas Guzmán et al., 2024), multi-step inference (Creswell et al., 2023), and consistency across formal reasoning patterns (Parmar et al., 2024; Hong et al., 2024). Neuro-symbolic methods address these gaps by integrating logic modules or symbolic solvers, improving both performance and interpretability (Pan et al., 2023; Olausson et al., 2023; Kambhampati et al., 2024). In a different direction, LRMs have shown strong gains in reasoning and planning tasks (Xu et al., 2025). Our proposed meta-learning approach offers a complementary alternative by enabling LLMs to adapt across logical tasks without relying on symbolic modules, as our results demonstrate.
@@ 8 Conclusion
In this work, we introduced MIND, a meta-learning fine-tuning approach to improve deductive reasoning in LLMs, explicitly targeting the logical premise selection task. Our results show that MIND significantly enhances generalization compared to the baseline, especially in small-scale and low-data scenarios. Remarkably, our fine-tuned small models outperform state-of-the-art LLMs on this task. This demonstrates the potential of MIND to advance the development of more robust and reliable AI systems.
Future work should explore several potential avenues. First, we should investigate not only systematic generalization using fixed inference rules, as we have done here, but also extend our research to learning the composition of multiple logical inferences. This approach aligns with ideas proposed in other meta-learning research, such as Meta-Learning for Compositionality(Lake and Baroni, 2023). Additionally, we should examine increasingly complex fragments of language, where the interactions among various inference-building blocks and reasoning forms become more intricate, and assess the effectiveness of MIND in helping LLMs to generalize in such contexts.

### v2

@@ 4 Experimental Setup
4.1 Models and Evaluation
We run experiments using the Qwen 2.5 family of decoder-only LLMs (Qwen Team, 2025). More specifically, we test three sizes: 1.5B, 3B, and 7B parameters. This family of models is selected because it allows us to experiment with varying small sizes (ranging from 1.5 to 7 billion parameters) and achieves a better size vs. performance trade-off compared to other open weights model families.
In addition to the Qwen 2.5 family, we also evaluate the closed-source LLM GPT-4o (OpenAI, 2024) and the Large Reasoning Model (LRM) o3-mini (OpenAI, 2025) on the logical premise selection task. [^8] Note that LRMs are also LLMs, but post-trained to generate longer intermediate chains of thought, improving performance on complex reasoning tasks (Xu et al., 2025). We evaluate these latter models both in a zero-shot setting and in a few-shot setting, using the $S^{\mathrm{supp}}$ study pairs as examples. [^9] See the API details and the exact prompts used to evaluate closed models in Appendix C.3.
@@ 4.1 Models and Evaluation
Fine-tuning
Qwen-2.5 1.5B ML 93.11 ± 0.61 94.28 ± 0.61 91.76 ± 0.27
$\mathrm{Baseline}_{-S}$ 85.56 ± 1.24 91.42 ± 0.82 80.56 ± 1.78
Qwen-2.5 3B ML 96.16 ± 0.44 96.24 ± 0.56 95.55 ± 0.43
$\mathrm{Baseline}_{-S}$ 93.03 ± 1.15 95.34 ± 1.18 90.92 ± 1.27
Qwen-2.5 7B ML 98.13 ± 0.98 98.26 ± 0.82 97.69 ± 1.40
$\mathrm{Baseline}_{-S}$ 95.76 ± 1.10 97.27 ± 1.22 94.13 ± 0.90
Prompting
GPT-4o Few-shot 39.76 52.91 33.51
@@
Table 1: Core generalization. Accuracy (mean ± std) on test inferences across all type‐length combinations (All), plus breakdown into the five shortest (Short) and longest (Long) inferences for each of the seven types of inference. Fine-tuned Qwen models use ML vs. Baseline; GPT-4o and o3-mini use few-shot vs. zero-shot prompting.
We design experiments to evaluate the ability of ML to teach pretrained small LMs to systematically apply inferences to new, unseen sets of premises —that is, to reason in a formal way by recognizing and instantiating the same underlying structure independently of the $\mathcal{KB}$s’ content.
@@
We then consider two more challenging generalizations involving inferences of unseen length. As illustrated in Figure 3, we examine the case of generalizing to longer inferences when the model has only learned from shorter ones (as studied in Saparov et al. 2023), and vice versa —generalizing to shorter inferences after seeing only longer ones. In the logic literature, they are respectively known as recursiveness and compositionality (Vargas Guzmán et al., 2024). To test this, we train exclusively on inferences whose lengths $x$ are $\sigma(t)\leq x\leq\mu(t)-5$, and test on the five longest inferences for each type, i.e., those whose length is $\mu(t)-5<x\leq\mu(t)$. In the second case, we train on inferences with length $\sigma(t)+5\leq x\leq\mu(t)$, and test only on the five shortest inference lengths for each type, i.e., those with length $\sigma(t)\leq x<\sigma(t)+5$. An intuitive representation of these generalizations is provided in Figure 3. Notably, we consider two varying types of study examples $S^{\mathrm{supp}}$: the aligned and disaligned sets of study examples, in which each $(x^{\mathrm{supp}},y^{\mathrm{supp}})$ either falls within or outside the range of inference lengths used for testing, respectively. [^10] More precisely, aligned and disaligned depend on whether we are evaluating models on unseen shorter or longer inferences. For longer inferences, disaligned includes inferences with lengths $\sigma(t)\leq x\leq\mu(t)-5$, and aligned includes those with lengths $\mu(t)-5<x\leq\mu(t)$. For shorter ones, instead, aligned includes inferences with lengths $\sigma(t)\leq x<\sigma(t)+5$, and disaligned includes those with lengths $\sigma(t)+5\leq x\leq\mu(t)$.
Figure 7, in the Appendix, shows all inference type-length combinations within training and test split in the core and in the length generalization settings. These datasets contain 1,000 and 100 datapoints for each training and testing type–length combination, respectively. To further investigate the performance of ML in a limited data regime, we also consider the case where only 100 datapoints are available for each training type–length combination.
@@ 4.2 Prediction Accuracy
We consider a model prediction to be correct if the set of premises extracted from the generated text matches the ground truth set of minimal premises. Using this criterion, we measure accuracy both aggregated, i.e., across an entire test set, and decomposed by each test type-length combination. All models (1.5B, 3B, and 7B) are fine-tuned three times and with different random seeds, thus we report mean and standard deviation of each accuracy.
@@ 5 Results
Qwen-2.5 1.5B ML 76.42 ± 2.95 91.75 ± 1.10 70.94 ± 2.27 71.13 ± 1.83
$\mathrm{Baseline}_{-S}$ 63.53 ± 1.16 63.53 ± 1.16 56.67 ± 1.22 56.67 ± 1.22
Qwen-2.5 3B ML 87.61 ± 1.97 95.86 ± 0.70 77.19 ± 3.53 78.53 ± 1.71
$\mathrm{Baseline}_{-S}$ 76.78 ± 1.63 76.78 ± 1.63 71.88 ± 1.49 71.88 ± 1.49
Qwen-2.5 7B ML 90.03 ± 1.09 96.84 ± 0.15 76.23 ± 2.91 83.41 ± 1.63
$\mathrm{Baseline}_{-S}$ 80.76 ± 2.65 80.76 ± 2.65 71.08 ± 1.55 71.08 ± 1.55
Table 2: Generalization to unseen lengths. Accuracy (mean ± std) of meta-learning and baseline models when trained on short inferences and tested on longer ones or vice versa. In both cases, we compare the settings in which the inferences in the study examples either falls within (Aligned) or outside (Disaligned) the range of inference lengths used for testing. $\mathrm{Baseline}_{-S}$ models have no study examples, hence such difference does not hold for them.
@@ 5.1 Core Generalization
Across all Qwen-2.5 model sizes (1.5B, 3B, 7B), the meta-learning approach consistently yields higher accuracy than the baseline. Performance improves with model scale in both approaches. For example, ML accuracy increases from 93.11% (1.5B) to 98.13% (7B) on all type‐length combinations, with accuracy on shortest inferences rising from 94.28% to 98.26%, and on the longest ones increasing from 91.76% to 97.69%. In contrast, baseline performance rises more slowly —from 85.56% (1.5B) to 95.76% (7B) —and shows a wider drop on the longest inferences, falling as low as 80.56% for the smallest model. Notably, the performance gap between ML and the baseline narrows as model size increases, suggesting that larger models achieve better core generalization even without meta-learning. It is worth noting that with limited data, ML’s advantage over the baseline becomes much wider at all sizes, as shown in Appendix D.3.
The closed-source models GPT-4o and o3-mini still underperform compared to Qwen-2.5 models fine-tuned with ML. Both models perform poorly in the zero-shot setting but improve with few-shot prompting: GPT-4o reaches 39.76% on all type‐length combinations (with 52.91% on shortest and 33.51% on longest inferences), while o3-mini performs substantially better (88.45% all combination, 87.91% on shorters, and 88.51% on longest). As expected, performance on the longest inferences is worse than that on the shortest ones for GPT-4o, while o3-mini maintains a more robust performance across inference lengths.
@@ 5.2 Length Generalization
Table 2 shows that ML models consistently outperform baseline models in generalizing to both longer and shorter inferences than those seen during training. In core generalization, we observed that longer inferences are more challenging than shorter ones. Instead, in the case of unseen lengths, an interesting and somewhat counterintuitive pattern emerges: it is generally easier for models to generalize to longer inferences than to shorter ones. This is true across all model sizes and in both approaches; for instance, the largest model, Qwen-2.5 7B, achieved 90.03% accuracy on longer inferences (disaligned) compared to 76.23% on shorter ones (disaligned).
Aligning study example lengths with the test condition (aligned) proves moderately to highly effective for unseen short inferences across all ML model sizes. For example, Qwen-2.5 1.5B improved from 76.42% to 91.75%, and Qwen-2.5 3B improved from 87.61% to 95.86%. For unseen long inferences, this alignment is moderately effective in larger models: Qwen-2.5 7B improved from 76.23% to 83.41%, while the 1.5B and 3B models showed smaller gains (70.94% to 71.13% and 77.19% to 78.53%, respectively). These results indicate that ML enables models in the aligned condition to exploit abstract patterns in the study examples (unseen inference lengths), allowing them to more effectively answer query hypotheses requiring length generalization.
Again, ML’s better performance in length generalization is especially noticeable with limited training data, where the difference between ML and baseline models grows significantly (see Appendix D.3 for more details).
@@
L $\rightarrow$ S ML (aligned) 42.94 4.9 36.68 2.1 57.5
ML (disaligned) 28.31 3.72 52.81 1.76 66.06
$\mathrm{Baseline}_{-S}$ 28.21 6.19 23.38 2.1 72.78
S $\rightarrow$ L ML (aligned) 9.76 1.66 87.54 5.08 60.94
ML (disaligned) 14.14 6.14 81.82 3.65 35.35
$\mathrm{Baseline}_{-S}$ 3.87 2.36 89.79 6.66 66.9
Table 3: Error analysis. Error analysis comparing ML and baseline on long to short (L $\rightarrow$ S) and short to long (S $\rightarrow$ L) generalization. The table shows percentages and averages for non-minimal valid sets of premises (NVM) and missing necessary $A$ premises (MAP), and the percentage of hallucinated premises (HP).
@@ 6 Error Analysis
Beyond simply measuring the accuracy of ML and the baseline, we additionally focus on two main types of errors models make when evaluated on unseen lengths. First, among all errors, we consider the proportion of non-minimal valid set of premises (NVM). This means that the correct minimal set was generated by the model, but together with unnecessary premises; for this case, we also measure how many unnecessary premises, on average, the models generate. Alternatively, models may fail to provide the complete $A$-chain within the correct minimal set of premises, meaning that at least one necessary $A$ premise is missing (MAP); here, we also track the average number of missing necessary $A$-formulas in erroneous answers. NVM and MAP are mutually exclusive. Furthermore, we consider an additional type of error that can occur simultaneously with either NVM or MAP: models may hallucinate premises —referred to as hallucinated premises (HP) —and output a formula that is not contained in the $\mathcal{KB}$.
Table 3 presents the error analysis for Qwen-2.5 7B [^11] Each model was fine-tuned three times with different random seeds, we selected the best model for each approach for this analysis. on the challenging length generalization settings. [^12] See Appendix D.5 for further error analysis results. HP is a common error type across both settings (often >>50%). The baseline model has the highest HP rate in long to short (72.78%), while ML models are generally better.
When generalizing to shorter inferences, a substantial portion of errors (28-43%) are NVM, indicating models indeed find logical solutions but include unnecessary premises. In this context, a lower number of unnecessary premises is better, as it is closer to the minimal set. The baseline model adds the most unnecessary premises (6.19 average), compared to ML (disaligned) (4.9) and ML (aligned) (3.72).
For generalizations to longer inferences, errors show different patterns, with few NVM errors (4-14%) and predominantly MAP errors (81-90%). The average number of missing premises is higher in short to long (3.65-6.66) than in long to short (1.76-2.1), suggesting models struggle to provide the complete set of premises when evaluated on longer inferences than seen during training. The baseline model struggles most with longer inferences, with a high MAP error rate (89.79%) and a large number of missing premises (6.66) contributing to its lower accuracy compared to ML.
@@ 7.1 LLMs’ Logical Capabilities
LLMs often struggle with OOD generalization (Clark et al., 2021; Saparov et al., 2023; Vargas Guzmán et al., 2024), multi-step inference (Creswell et al., 2023), and consistency across formal reasoning patterns (Parmar et al., 2024; Hong et al., 2024). Neuro-symbolic methods address these gaps by integrating logic modules or symbolic solvers, improving both performance and interpretability (Pan et al., 2023; Olausson et al., 2023; Kambhampati et al., 2024). In a different direction, LRMs have shown strong gains in reasoning and planning tasks (Xu et al., 2025). Our proposed meta-learning approach offers a complementary alternative by enabling LLMs to systematically learn logical relations.
@@ 8 Conclusion
In this work, we have shown that ML is a promising approach to enhance the logical capabilities of small LMs, at least with respect to the syllogistic fragment. We applied few-shot meta-learning as a fine-tuning approach to improve the deductive reasoning of models, explicitly targeting the logical premise selection task. Our results show that ML significantly enhances logical generalization compared to the baseline, especially in small-scale and low-data scenarios. Remarkably, our fine-tuned small models outperform state-of-the-art LLMs on our syllogistic reasoning task. This demonstrates the potential of ML to advance the development of more robust LLM reasoners.
In the future, we should investigate not only systematic generalization using fixed inference rules, as we have done here, but also extend our research to learning the composition of multiple logical inferences. This approach aligns with ideas proposed in other meta-learning research, such as Meta-Learning for Compositionality (Lake and Baroni, 2023). Additionally, we should examine increasingly complex formal fragments of language, where the interactions among various inference-building blocks and reasoning forms become more intricate, and assess the effectiveness of ML in helping LLMs to generalize in such contexts. Furthermore, having shown that in principle ML may enhance logical generalization, we want to next study the effectiveness of ML in more naturalistic or real-world reasoning scenarios.